OpenStudy (anonymous):

write 6 4/7 as an improper fraction

OpenStudy (igreen):

Multiply the denominator to the whole number, then add it to the numerator.

OpenStudy (igreen):

Can you do that? @Tttttttcoockiemonster

OpenStudy (igreen):

\(\dfrac{1\leftarrow Numerator}{2 \leftarrow Denominator}\)

OpenStudy (ahsome):

Improper fraction occurs when the numerator is bigger than the denominator Your equation is: \[6\frac{4}{7}\] Multiply the denominator by the whole number? \[6*7=42\] Use that for the numberaot, plus the 4 \[\frac{42+4}{7}\] \[\frac{44}{7}\]

OpenStudy (anonymous):

46/7

OpenStudy (ahsome):

46, Yeah

OpenStudy (igreen):

Yeah xD

OpenStudy (igreen):

Great job @Tttttttcoockiemonster